# Plan2Align: Predictive Planning Based Test-Time Preference Alignment in Paragraph-Level Machine Translation

### v1

@@ 3.1.1 Context Alignment and Context Buffer
Each translation is aligned through context alignment, establishing correspondences between source and translation contexts. We evaluate the aligned source and translation context pairs using a reward model trained on given preference data. Our goal is to ensure that the context buffer retains only high-quality contexts. The reward model assesses candidate context-level translations and selectively retains only the best ones. When a selected translation context surpasses a predefined buffer threshold and a matching source context already exists in the context buffer, a comparison is performed. If the new context falls below the threshold or does not surpass the quality of the existing context in the buffer, it is not stored.
Through this approach, in each iteration, the system filters high-quality contexts, continuously refining the context buffer and ensuring that subsequent translations improve over time.
3.1.2 Translations from Self-Rewriting Tasks
To leverage high-quality contexts in the context buffer for translation improvement, we frame the translation task as a self-rewriting task. To increase the diversity of translation outputs generated by the LLM, thereby enriching the context buffer with more varied contexts, we sample reference contexts from the buffer to construct the self-rewriting task input, rather than using the entire buffer content. By providing LLMs with pre-aligned content as input for self-rewriting, they can refine paragraph-level translations using past high-quality contexts while maintaining context-level coherence.
3.2 Final Translation from Context Buffer
The final translation generation differs from context buffer updates. In this phase, the stored high-quality contexts, covering nearly the entire paragraph, serve as reference translations. Unlike the self-rewriting task, which aligns individual sentences, this final translation generation stage constructs the reference translation by incorporating all relevant content from the context buffer and feeding it along with the source paragraph into the LLM. The LLM then improves the reference translation based on the source, preventing overlapping translation artifacts and ensuring fluency. Since this step involves refining a single high-quality reference translation with the source text, even non-extensive LLMs can handle the task effectively.
3.3 The Details of Context Alignment
Since the final translation relies on the quality of the context buffer, improving its quality is crucial. Therefore, continuously updating the context buffer to retain the most valuable translation experience is essential. To achieve this, context alignment must ensure proper pairing between source and translation contexts, allowing the reward model to effectively filter out low-quality translations and retain only the most reliable ones.
However, there is no well-established method to effectively extract contextual information in paragraph-level MT. To address this, we propose a simple yet effective approach for context alignment, which consists of three key steps: (i) Sentence Segmentation: We utilize the SpaCy (Honnibal et al., 2020) toolkit to segment paragraphs into a list of sentences. (ii) Sentence Alignment: We use VecAlign (Thompson and Koehn, 2020) to dynamically search for the optimal alignment between sentences. Since LLMs may introduce hallucinations, leading to sentence omissions or over-translations (resulting in extra sentences), we insert any unaligned translation sentences into their original positions and add empty strings on the source side as placeholders. (iii) Contextual Sliding Window Construction: To capture the contextual structure at the paragraph level, we adopt a sliding window approach with a window size of 3, forming the context-level aligned source-translation pairs.
@@ 3.3 The Details of Context Alignment
[…]
Llama-3.1Vanlia ${}_{\text{Vanlia MPC}}$ 89.79 58.40 7.30 4.55 82.83 35.18 7.19 10.26 83.67 38.46 10.82 10.63
@@
4 Experiments
4.1 Translation Dataset
Compared to sentence-level MT, paragraph-level MT has received significantly less attention, and many translation models struggle to process long-context inputs. For example, Post and Junczys-Dowmunt (2024) investigated a long-context training paradigm for machine translation but restrict the maximum sequence length to 256. ALMA Xu et al. (2024a, b) restricted their post-training datasets to a maximum sequence length of 512. To address this limitation and ensure a suitable benchmark for paragraph-level MT, we use the latest WMT’24 Discourse-Level Literary Translation benchmark (Wang et al., 2024a) for our experiments. This dataset allows for evaluating how well translation methods preserve contextual relationships during translation. The available language pairs include: Chinese →English, Chinese →German, and Chinese →Russian.
To ensure that the input length remains within the context window of our selected LLMs, we segment each instance from the training and validation sets into chunks of up to 1024 tokens. This preprocessing ensures that paragraph-level MT tasks are effectively handled within the model’s constraints. But as the readers shall see, our methodology is easy to extend to even longer context given LLM translation capability is up to par.
4.2 Preference Data and Model Training
Preference data is crucial for training alignment methods and for training the reward model used in Plan2Align. The preference data is derived from the training set of the dataset. Specifically, each instance is segmented into paragraphs of up to 1024 tokens. From each translation direction, we sample 2K paragraphs, resulting in a total of 6K paragraphs for constructing the preference data. We generate translation outputs using LLaMA-3.1-8B-Instruct, Gemma-2-9B, and GPT-4o. The translations are then evaluated using MetricX-24-XL (Juraska et al., 2024), a reference-free quality estimation (QE) model. The QE scores help us select high-quality translations: the translation with the highest score is designated as the preferred translation, while the one with the lowest score is considered not preferred. The translation with a middle score is disregarded. The resulting preference data is used to train the reward model in Plan2Align, enabling it to assess context quality. The trained reward model achieves a final accuracy of 88.53% on the validation set. Further details on the formation of preference data can be found in Appendix A.
4.3 Evaluation Metrics
We focus on optimizing LLM-based paragraph-level MT models to mitigate contextual inconsistencies, hallucinations, and omissions. However, while model-based metrics have taken over string-based metrics as the state-of-the-art, all model-based metrics (e.g. COMET) are still only trained with sentence-level data, and machine translation evaluation on paragraph-level remains an open question. A prior work Deutsch et al. (2023) argued that existing metrics generalize well to paragraph-level, but the test set in that work is a mere concatenation of existing datasets translated by sentence-level models, and does not take into account the context-level improvements we are aiming at. In fact, in our preliminary study, we observed serious pathological behaviors of these metrics when evaluating our paragraph-level outputs (see Appendix E). This calls for a significant rework of the existing evaluation paradigm.
The changes we’ve made to our machine translation evaluation paradigm are as follows:
Contextual Sliding Window.
Instead of simply feeding the source, translation and reference sentences into COMET [^1] Unbabel/wmt22-comet-da and COMET-KIWI (the reference-free version) [^2] Unbabel/wmt22-cometkiwi-da, we follow Vernikos et al. (2022) and add aligned and concatenated context as inputs to the model-based metrics. We call these metrics CW-COMET and CW-KIWI. We use the same context size of 3 as described in Section 3.3.
Sentence Segmentation and Alignment.
Similar to IWSLT evaluation paradigm for speech translation Ahmad et al. (2024), we propose to treat sentence segmentation and alignment as part of the evaluation for our paragraph-level evaluation. We follow the same general procedure as described in Section 3.3. We handle sentence alignment differently according to whether the metric requires reference input. Misaligned sentences remain in their original positions, with empty strings as placeholders where necessary. If a sliding window contains only empty strings on the one side, a default score of zero is given. This penalizes over-translation and omission, aligning with our research objectives. We verified the validity of our paradigm via a sanity check. Details are in Appendix D.
In addition to reporting model-based metrics, because over- and under-translations usually create significant misalignments across sentences, we track them by reporting the ratio of 1-0 and 0-1 alignments (1-0/0-1 Ratio), which represents the proportion of windows in the source or translation that fail to align. This is evaluated on our sentence-segmented and sentence-aligned test set. We also report document-level BLEU (d-BLEU) as some previous work has done, even it does not correlate well with human judgments Deutsch et al. (2023) and should only be treated as an auxiliary metric.
4.4 Baselines
We evaluate all training-time alignment methods on LLaMA-3.1-8B-Instruct and also use it as the backbone model for test-time alignment methods, including Plan2Align. This ensures consistency in comparing the effectiveness and aligns with our focus on small-scale LLMs. Additionally, we compare these results against the original LLaMA-3.1-8B-Instruct to further assess improvements and changes, serving as the baseline for performance. Implementation details of Plan2Align, including parameters and prompt design, are in Appendix C.
GPT-4o and Qwen-2.5-14B.
We select GPT-4o as the upper bound for performance. GPT-4o is not optimized for translation, but closely approaches that of specialized translation systems (Shahriar et al., 2024). Its precedent model, GPT-4, was the top-ranked model for 5 langauge pairs in WMT24 general machine translation shared task Kocmi et al. (2024). In addition, we also include Qwen-2.5-14B for comparison, as it demonstrates strong performance in Chinese language contexts.
Test-Time Alignment Models.
We select ARGS (Khanov et al., 2024) and RAIN (Li et al., 2024) as they both represent test-time alignment methods similar to Plan2Align but with distinct operational mechanisms. ARGS is selected for its reward-guiding method similar to Plan2Align’s; however, while Plan2Align evaluates the entire context with a reward model, ARGS dynamically recalculates scores for each token during generation. On the other hand, RAIN is selected for its ability to score newly generated tokens and perform auto-regressive self-evaluation, requiring a start-from-scratch search and backward progression. To better observe the benefits brought by MPC, we also compare it with Best-of-N and Vanilla MPC. The implementation details of test-time alignments, are available in Appendix B.1.
Training-Time Alignment Models.
We also compare Plan2Align with various training-time alignment methods. We consider SFT directly on the same preference dataset. Moreover, we include SimPO for comparison as it serves as a strong benchmark for training-time alignment in general. We also consider DPO as it serves as a mainstream representative of the training-time-alignment series, with training details available in Appendix B.2.
4.5 Quantitative Results
We present the primary results for zh→xx in Table 1. Our primary goal is not to outperform all models but rather to maximize the utility of existing models, approaching or even surpassing training-time alignment methods. Since training-time methods have direct exposure to real paragraph data, direct comparisons to test-time methods would be unfair. Instead, we emphasize whether Plan2Align can consistently improve existing model performance. We summarize the observations as follows:
Plan2Align Outperforms Existing Test-Time Alignment Methods.
Table 1 demonstrates that Plan2Align outperforms test-time alignment baselines across all translation directions. This is primarily because paragraph-level MT requires generating long outputs, and existing methods such as ARGS and RAIN guide the decoding process at the token level through a reward model. These methods struggle with error accumulation, where poor early-context translations negatively affect subsequent generations. Additionally, Table 2 highlights the efficiency of Plan2Align compared to ARGS and RAIN, detailing the time required for each approach across iterations. Since Plan2Align operates directly at the context level, it generates complete translated paragraphs without additional decoding overhead, making it both faster and more effective.
Plan2Align Closes the Performance Gap Between Test-Time and Training-Time Alignment Methods.
[…]
Best-of-N, Vanilla MPC, and Plan2Align.
As discussed in Section 1, non-extensive LLMs struggle to generate high-quality paragraph translations in a single pass. This limitation renders Best-of-N ineffective for paragraph-level MT, while directly applying MPC (i.e., Vanilla MPC) fails due to the model’s difficulty in selecting optimal contexts for subsequent translations. Table 1 shows that Best-of-N improves translation quality but does not enhance the model’s fundamental capabilities. Its effectiveness remains limited by randomness, making it an inefficient approach. Even with $N=60$ (six times the number of iterations in Plan2Align), it still underperforms compared to Plan2Align across all translation directions. Vanilla MPC achieves improvements for zh→en, but its effectiveness diminishes in zh→ru and zh→de, sometimes even falling below LLaMA-3.1’s baseline performance. This suggests that for less familiar languages, LLMs struggle to extract meaningful improvements. Plan2Align overcomes this limitation through context alignment and self-rewriting, enabling the effective application of MPC to paragraph-level MT.
AVG. Time Consumed
[…]
ARGS 1686 2343 2035
RAIN 4639 5518 4454
Plan2Align 311 313 424
Table 2: Average execution time (in seconds) per paragraph for three test-time alignment methods on the validation set. Plan2Align, with an iteration number of 3, is significantly faster than ARGS and RAIN since it avoids token-level alignment. Tests are conducted on a single NVIDIA RTX 6000 Ada Generation GPU.
[…]
Plan2Align 1 93.28 66.89 20.99 0.32
2 93.82 66.72 20.75 0.52
3 94.13 67.06 21.17 0.20
4 93.95 66.19 20.81 0.38
5 93.97 66.59 21.09 0.35
Vanilla MPC 1 89.79 58.40 17.30 4.55
2 88.00 56.26 16.45 6.42
3 83.92 53.08 15.11 10.70
4 84.66 53.60 15.26 9.93
5 84.93 54.03 15.48 9.68
Table 3: Performance of Plan2Align and Vanilla MPC across different iterations on the zh $\rightarrow$ en task.
Impact of Iteration Number in MPC.
As iterations increase, the context buffer expands, and Plan2Align selects more reference contexts. However, over-translation and omission often prevent perfect semantic alignment. Using excessive reference contexts in the self-rewriting task can overload the LLM, making it harder to extract meaningful insights due to overlapping semantics. To isolate translation difficulty effects, we focus on zh $\rightarrow$ en performance. Table 3 shows that while increasing reference contexts improves translation initially, excessive references (iteration = 4) degrade performance due to semantic ambiguity. Interestingly, Vanilla MPC exhibits an even sharper decline that is starting at iteration = 2, even in the simpler zh $\rightarrow$ en task. This supports our hypothesis that LLMs cannot directly leverage MPC effectively. Small-scale models struggle to extract useful contexts, leading to error accumulation and deteriorating translations. Plan2Align overcomes this limitation through structured context alignment and self-rewriting, ensuring stable improvements in paragraph-level MT.
5 Conclusion
In this paper, we introduced Plan2Align, a new form of model predictive planning framework for test-time preference alignment in paragraph-level machine translation. By recasting translation as a trajectory optimization problem and leveraging a self-rewriting mechanism, our approach effectively addresses key challenges such as semantic drift, omissions, and incoherence that arise in one-pass, next-token generation. Our results suggest that predictive planning can robustly accumulate and utilize high-quality translation contexts, paving the way for further improvements in long-form machine translation. Future work will explore broader language pairs and the integration of training-time strategies to further enhance translation quality. Our code will be open source under MIT license.
Limitation
One of our limitation lies in the scope of translation languages. Our experiments have focused on language pairs (e.g., Chinese to English, German, and Russian) from reproducible benchmarks, and while these choices provide valuable insights, they could potentially not encompass the full linguistic diversity present in the other translation tasks. Languages with complex morphologies, low-resource languages, or those with radically different syntactic structures may present challenges that our current formulation of context alignment and iterative refinement does not fully address. Future work has been scheduled to explore a broader set of languages to ensure that the method generalizes well and does not inadvertently favor certain language families or linguistic features over others.
Ethical and Societal Considerations
While Plan2Align is designed to enhance translation coherence and quality, its deployment raises important ethical and societal considerations. First, there is an inherent risk of bias amplification. The reward model, which drives context selection during test-time alignment, is trained on preference data that may itself contain implicit biases. If not carefully audited and diversified, this could lead to translations that systematically favor dominant linguistic norms or cultural perspectives, marginalizing underrepresented human dialects. In sum, by focusing on a limited set of languages and model architectures, there is a risk of reinforcing existing digital divides. If high-quality translation tools are predominantly developed for well-resourced languages and large-scale models, speakers of low-resource or underrepresented languages may be left behind. As researchers and practitioners, it is a long-term developing responsibility to prioritize inclusivity and fairness, ensuring that technological advances in machine translation benefit a diverse research community.
Acknowledgment
This research has been supported by NVIDIA grant and technical involvement from NVIDIA Taiwan AI R&D Center (TRDC).
Appendix A Details of Preference Data
We used MetricX-24-XL to labeled our pairwise preference datasets, the Chinese sources sentences are from WMT’24 Discourse-Level Literary Translation benchmark, we employed LLaMA-3.1-8B-Instruct, Gemma-2-9B, and GPT-4o to generate translations across three language pairs; each language pair has 2000 records, with a maximum length of 1024 tokens. The distribution of preferences, indicating the number of translation is the best translation among the three,indicating the number of translation is the best translation among the three, given that MetricX-24 scores range from 0 to 25, where 0 is the best and 25 is the worst. We removed translations scoring above 20, and if two out of three translations in a paragraph exceeded this threshold, we did not use that paragraph. The details are in Table 4. Our test-time setting, pre-trained models, and preference data will be open source under MIT license.
LLaMA Wins Gemma Wins GPT4 Wins
[…]
Appendix B Implementation Details of Baselines
B.1 Test-Time Alignment Methods
B.1.1 ARGS

### v2

@@ 4.4 Baselines
Llama-3.1 (8B) $\times$ 2.9528 0.6160 0.7936
[…]
Llama-3.1Vanlia ${}_{\text{Vanlia MPC}}$ 4.0776 0.6364 0.7885
[…]
Figure 3: Results on the test set of Dahoas/full-hh-rlhf. Left: Quantitative comparison across base models, training-time, and test-time alignment methods. For the two groups showing results from alignment methods, best numbers in respective groups are boldfaced, and best numbers among all alignment methods are also underlined. Our method are highlighted. Right: Win rate of Plan2Align responses compared to DPO and Best-of-20.
4.5 Quantitative Results
4.5.1 Results on Long-Form Responses
We present the primary results for long-form response generation in Figure 3. This task provides a challenging benchmark for evaluating how well different alignment methods can enhance the generation quality of base model. We summarize the observations as follows:
Superior to training-time and test-time baselines.
Compared to strong baselines such as DPO and SimPO, our method achieves higher average reward, coherence, and diversity, all without any model fine-tuning. Moreover, it delivers a substantial improvement over the base model. We further assess response quality using GPT-4 preference judgments on the 1,024 test samples. The reward model trained with Plan2Align achieves an accuracy of 75.71% on this evaluation set, demonstrating strong alignment with GPT-4 preferences. To provide a comparative analysis, we evaluate Plan2Align against the best-performing methods from both training-time and test-time alignment: DPO and Best-of-20, respectively. As illustrated in the right side of Figure 3, Plan2Align significantly outperforms both baselines in head-to-head comparisons. GPT-4’s win/loss/tie statistics for each comparison are detailed in the figure
Effective alignment with fewer rollouts.
Our method performs 3 iterations with 3 rollouts each, in addition to the initial LLM output, totaling 10 generations. In contrast, Best-of-20 samples twice as many outputs, yet still underperforms relative to Plan2Align. This demonstrates that our gains are not attributable to random sampling. Even under equal generation cost (Best-of-10), Plan2Align maintains a clear advantage, validating the benefit of planning and rewriting over naive sampling. Despite its improved alignment quality, Plan2Align remains efficient at inference time. While the baseline LLaMA3.1-8B model requires 8 seconds to generate a single response, ARGS takes 363 seconds (44×), and RAIN takes 1930 seconds (242×). In contrast, Plan2Align completes inference in 108 seconds—only 13× the cost of vanilla generation—while being substantially more efficient than other test-time alignment methods. These results highlight that Plan2Align achieves a strong balance between alignment performance and generation efficiency.
Ablation via Vanilla MPC.
In this variant, we remove segment-level evaluation and the reference buffer, selecting only the best full response at each iteration for refinement. While this strategy still improves upon the base model, it lags behind both Plan2Align and Best-of-10, highlighting the necessity of our core design choices for effective preference alignment.
zh $\rightarrow$ en zh $\rightarrow$ ru zh $\rightarrow$ de
[…]
GPT-4o 2024-08-06 ${}_{\text{~{}2024-08-06}}$ 94.58 73.06 93.74 54.20 94.54 54.55
Qwen-2.5 (14B) - 94.43 72.44 90.47 50.13 92.98 50.02
Llama-3.1 (8B) $\times$ 84.36 58.27 86.28 34.32 88.97 39.49
[…]
Llama-3.1Vanlia ${}_{\text{Vanlia MPC}}$ 88.50 57.06 73.84 31.45 90.32 42.20
[…]
4.5.2 Results on Paragraph-Level MT
We present the primary results for zh→xx in Table 1. Compared to HH-RLHF, we consider long-text translation a more challenging task, especially for small-scale LLMs. This is reflected in Table 1, where we observe a large gap between the base model and SoTA models. Therefore, our primary goal is not to outperform all models but rather to maximize the utility of existing models, approaching or even surpassing training-time alignment methods. We summarize the observations as follows:
Plan2Align Narrows the Gap Between Test-Time and Training-Time Alignment.
As shown in Table 1, Plan2Align consistently outperforms all existing test-time alignment baselines across translation directions. Unlike token-level methods such as ARGS and RAIN, which guide decoding with fine-grained reward signals and suffer from early error accumulation, Plan2Align operates at the segment level and incorporates high-quality partial translation into a rewriting framework. This enables greater coherence and contextual consistency in paragraph-level translation. Moreover, Plan2Align narrows the gap with training-time alignment methods, despite lacking the direct exposure to paragraph-level data that fine-tuning approaches typically benefit from. For zh→en, Plan2Align outperforms all training-time baselines under the CW-COMET metric, which is generally considered more reliable due to its reference-based nature. It even surpasses the GPT-4o and Qwen-2.5, highlighting the effectiveness of test-time planning in aligning long-form translations with human preferences. For zh→ru, although CW-COMET shows Plan2Align achieving strong performance, CW-KIWI scores remain lower than those of training-time methods. This discrepancy is likely due to limitations in the KIWI-22 evaluation model for this language pair, which fails to fully capture Plan2Align’s improvements. For zh→de, Plan2Align does not outperform training-time approaches, but it delivers comparable results. This suggests that while our method is generally effective, its performance may still be limited by the base model’s inherent translation abilities.
Best-of-N, Vanilla MPC, and Plan2Align.
Table 1 shows that Best-of-N sampling remains limited in effectiveness due to its reliance on unguided randomness. Even with $N=60$ (5 times the number of rollouts in Plan2Align), it still underperforms compared to Plan2Align across all translation directions. Similarly, the Vanilla MPC variant also fails to match the performance of Plan2Align. Without segment-level evaluation or rewriting, it lacks the structural guidance necessary for effective refinement. These results further validate the design of Plan2Align and its ability to perform efficient and robust preference alignment across domains.
5 Conclusion, Limitations and Future Work
We introduced Plan2Align, a test-time model predictive planning framework for preference alignment. By treating generation as trajectory optimization, Plan2Align incrementally refines outputs through segment-level evaluation and rewriting, leading to consistent improvements without modifying model parameters. While effective, Plan2Align depends on the base model’s expressiveness—it can only amplify behaviors already latent in the model. When desirable outputs lie far from the model’s distribution, test-time planning alone may be insufficient. Future work will explore how iterative distribution shifts arise under Plan2Align and how these mechanisms can be integrated with training-time alignment to support more robust generalization.
Appendix A Limitations and Future Work
While Plan2Align effectively improves output quality through test-time preference alignment, it is fundamentally constrained by the expressiveness of the underlying language model. It can only steer generation toward outputs already supported by the model’s capabilities, and may struggle when the desired outputs lie far from the model’s original distribution. In future work, we aim to better understand the conditions under which iterative alignment shifts the output distribution effectively. This could enable stronger integration with training-time methods, such as lightweight fine-tuning or reward model co-optimization, to improve robustness in harder settings and under distribution shifts.
Appendix B Broader Impacts
We anticipate that Plan2Align will contribute positively to the development of more controllable and aligned language models, particularly in low-resource or safety-critical applications where fine-tuning is impractical. By enabling preference-aware generation at test time without modifying model weights, our method may lower the barrier to deploying LLMs in real-world scenarios. However, test-time alignment methods also introduce new risks, such as reinforcing spurious patterns if reward signals are poorly calibrated. To mitigate this, we recommend careful design and auditing of reward models, especially in domains involving sensitive or subjective outputs.
Appendix C Details of Preference Data
Both of two reward model serves as the alignment objective for the Plan2Align framework.
C.1 Long-Form Responses
For the long-form response task, we construct a pairwise preference dataset using the Dahoas/full-hh-rlhf dataset, where each pair consists of a chosen and a rejected response. These pairs are used as preference data to train the reward model. To ensure a fair evaluation and maintain consistency in response length, we filter both the training and test sets to include only samples with continuation lengths close to 1024 tokens. Specifically, we select 6,000 training samples that fall within this length range.
C.2 Paragraph-Level MT
For paragraph-Level machine translation task, we used MetricX-24-XL to labeled our pairwise preference datasets, the Chinese sources sentences are from WMT’24 Discourse-Level Literary Translation benchmark, we employed LLaMA-3.1-8B-Instruct, Gemma-2-9B, and GPT-4o to generate translations across three language pairs; each language pair has 2000 records, with a maximum length of 1024 tokens. The distribution of preferences, indicating the number of translation is the best translation among the three,indicating the number of translation is the best translation among the three, given that MetricX-24 scores range from 0 to 25, where 0 is the best and 25 is the worst. We removed translations scoring above 20, and if two out of three translations in a paragraph exceeded this threshold, we did not use that paragraph. The details are in Table 2. Our test-time setting, pre-trained models, and preference data will be open source under MIT license.
LLaMA Wins Gemma Wins GPT4 Wins
[…]
Appendix D Supplementary Results
Since the results in the main paper are condensed versions with showing only a subset of the metrics evaluated, we attach the full breakdown of the results in Table 3 for readers’ reference. In addition to reporting model-based metrics, because over- and under-translations usually create significant misalignments across sentences, we track them by reporting the ratio of 1-0 and 0-1 alignments (1-0/0-1 Ratio), which represents the proportion of windows in the source or translation that fail to align. We also report document-level BLEU (d-BLEU) as some previous work has done, even it does not correlate well with human judgments (7) and should only be treated as an auxiliary metric.
@@ Appendix D Supplementary Results
Llama-3.1Vanlia ${}_{\text{Vanlia MPC}}$ 88.50 57.06 13.39 6.04 73.84 31.45 3.96 19.55 90.32 42.20 10.19 2.91
@@
Appendix E Evaluation prompt of GPT-4
Figure 4 is the prompt template for the GPT-4 evaluation, we follow the ARGS evaluation setup (15), with a slight modification:Our prompt explicitly instructs GPT-4 to consider how well each assistant addresses the user’s question, rather than to evaluate the response in isolation. The prompt asks GPT-4 to rate responses based on overall performance, considering criteria such as helpfulness, harmlessness, relevance, accuracy, depth, creativity, and level of detail
Figure 4: System prompt and user prompt of GPT-4 evaluation.
Appendix F Implementation Details of Baselines
F.1 Test-Time Alignment Methods
F.1.1 ARGS